OpenStudy (anonymous):

Solve the inequality. |r + 3| ≥ 7

OpenStudy (anonymous):

abs(r+3)>=7 -7<=r+3<=7

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

np

Directrix (directrix):

|r + 3| ≥ 7 comes apart as follows: (r + 3) ≥ 7 OR (r + 3) ≤ - 7 r ≥ 7 - 3 OR r ≤ -10 ------------------------- That is NOT the same as this from above: -7<=r+3<=7 @saraizzle

OpenStudy (anonymous):

oh you have a point, thanks :-)

OpenStudy (anonymous):

can you help me with this? Find the solutions of the inequality. | x - 5 | ≥ 3

Directrix (directrix):

If you submit this --> -7<=r+3<=7 as "the" answer, you will not get credit.